# Tracking Patterns in Toxicity and Antisocial Behavior Over User Lifetimes on Large Social Media Platforms

### v1

@@ Abstract
An increasing amount of attention has been devoted to the problem of “toxic” or antisocial behavior on social media. In this paper we analyze such behavior at very large scales: we analyze toxicity over a 14-year time span on nearly 500 million comments from Reddit and Wikipedia, grounded in two different proxies for toxicity.
@@ Introduction
Antisocial behavior on social media is recognized as an increasingly important problem for the experience of users on many large platforms. It has been implicated in patterns of online harassment,[1, 2] outcomes of civic processes,[3, 4, 5, 6] and physical violence.[7, 8] However, its underlying mechanisms are still poorly understood. One reason for this challenge is that the culture and dynamics of social media are also constantly changing, both within and across platforms.[9, 10] Users also influence and are influenced by the communities they interact with. [11, 12]
The present work is interested in whether patterns of toxicity change over time, not just globally across a platform, but particularly within individual users. We analyze trajectories across four dimensions: activity level (lifetime number of comments), two different metrics of antisocial behavior (Perspective API toxicity score and “downvote score”, discussed below), two different social media sites (Reddit and Wikipedia), and two types of histories (people/users and information/URLs).
In nearly all cases, we observe a striking reversal: users do change over their lifetimes, but the direction of that change hasn’t been stable over time. Before 2013, the longer a user spent on a social media site, the less toxic they became. After 2013, the longer they spent on the site, the more toxic they became. [^1] This pattern appears for both metrics used, but for the downvote metric, the reversal happens slightly later - around 2015.
We also observe another pattern: First, low-activity users are generally separable from higher-activity users in their toxicity behavior. On Reddit, higher-activity users have higher toxicity scores; on Wikipedia, the reverse is true. This raises the question of how this difference comes about: one possibility is that high-activity users start out looking exactly like low-activity users and gradually reach their final level of toxicity; a second possibility is that they have different toxicity from the beginning. We find that the data is more consistent with the latter principle: Low- and high-activity users have different average toxicity metrics from their very first post. That is, it doesn’t seem that users start out the same, then only become more or less toxic the more they comment. Instead, they’re distinct from each other from the beginning.
@@
We note at the outset that it’s worth being critical of the term “toxicity”. The term is commonly assumed to need no definition,[13, 14] and often used without explanation. But some think it’s too vague to be meaningful, or even that it’s used to avoid responsibility or to avoid more uncomfortable topics such as radicalization or identity-based hatred.[13, 15] It is therefore important to discuss the sense in which we will use it in our work.
@@
Google’s Jigsaw, the makers of the Perspective API,[16, 17, 18] give a practical reason for choosing the term: when training humans to annotate their training data for generally problematic text, they found that “toxic” was the easiest concept to understand and find an agreed-on definition for. That is, they tried various terms for various forms of problematic behavior, and the annotators were more likely to disagree on which comments met the definitions of any of the other terms.[19] For the API metric used in this work, “severe toxicity”, the specific definition they used was: “a hateful, aggressive, disrespectful comment or otherwise very likely to make a user leave a discussion or give up on sharing their perspective”.[20]
Whether or not we agree with their definition,[21, 22, 23] the model is clearly measuring some signal. And since we’re interested in patterns at large scale over long time periods, we think it can give us useful information. Plus, the fact that it’s commonly used in research[23] makes it easier to put results in context, and its use in real-world applications[24, 25] means that it’s worth studying.
@@
Another way to address the subtlety of the underlying concepts is to quantify them in multiple disjoint ways. Accordingly, we also supplement the Perspective API score with a completely unrelated metric: a “downvote score”, which also measures some kinds of antisocial behavior. Reddit’s content ranking depends on user voting, where any user can “upvote” or “downvote” a post or comment. Reddit asks users to vote based on whether a comment “contributes to conversation”, and is on-topic to the community it’s posted in.[26] (There’s no enforcement, however, and users sometimes vote based on other factors, such as whether they agree with the comment.) A comment’s score is a direct measure of how much its audience thinks it should be seen by others. Comments with a negative score, meaning they got more downvotes than upvotes, were likely judged in the eyes of their community to be antisocial in some way. As we will see, the broad conclusions we draw hold under both numerical metrics, despite the fact that they are based on essentially disjoint sources of information in the underlying data.
Many works look at online toxicity and antisocial behavior, many specifically using the Perspective API toxicity score.[27, 28, 29, 30] Other metrics of “toxicity” or “antisocial behavior” include banned users,[31] hand-coded “civility”,[32, 33] and alternate toxicity or hate speech classifiers.[34, 35] However, most of these are interested in conversation-level[30, 27, 35, 36, 32, 37, 38] or community-level[39, 40, 41, 42, 29] patterns.
@@
Of the few that look at individual trajectories over time, Cheng et al[31] analyze users who will go on to be banned from a community. Similar to this work, they find that “future banned users” are different from “never-banned users” from their very first post, and that these users’ behavior changes over time according to their main metrics: post deletion rate and text quality. They also measure slopes in users’ post deletion rates over time, though focus on whether the slopes change between the first and second half of a user’s lifetime. Almerekhi et al[34] follow user change at a more local level, using a binary definition of “toxicity” and counting events where a user made a “toxic” comment after a “non-toxic” comment or vice versa. Waller and Anderson[39] also find a temporal shift in other characteristics at a similar time to the shift found in this work. They found a sharp change in user polarization, measured by the partisanship of communities that users commented in. However, this polarization event wasn’t relative to user lifetimes; existing users all polarized at approximately the same time, though most of the polarization was driven by new users.
@@ Results
The toxicity score is from the Perspective API,[43] a machine learning model that assigns a piece of text a score between 0 and 1, where 1 is most toxic. The API defines “toxicity” as “a very hateful, aggressive, disrespectful comment or otherwise very likely to make a user leave a discussion or give up on sharing their perspective”.[20]
@@
The downvote score comes from the Reddit voting system, where any user can “upvote” or “downvote” a comment based on whether it “contributes to conversation”.[26] Here “downvote score” for a group of comments denotes the percentage of comments that were downvoted, meaning they received more downvotes than upvotes.
@@
The Reddit user dataset contains all comments made by a random sample of 1 million users ($\sim 1\%$), for a total of 123 million comments. The Wikipedia dataset contains all edits made to Wikipedia talk pages, after filtering out bot accounts, for a total of 122 million edits from 2.8 million accounts, on 21.8 million talk pages. Supplementary Section S.G shows the number of accounts made in different years.
The Reddit URL dataset consists of all Reddit posts of YouTube or Imgur URLs. Scores for each post were calculated as the mean of scores for all comments on the post. Downvote score for each post was defined as the percent of comments that were downvoted.
@@
As a first question, we consider whether users who comment often are different from more casual users in the metrics we study (Fig 1). On average, they are indeed different, with the most prolific users showing the highest average for both toxicity and downvote scores. For toxicity scores, there is a clear upward trend as we move towards more active users.
Black bars represent 95% confidence intervals from 1,000 bootstrap runs.
(To note, there are a number of options for stratifying users by activity. We use the lifetime number of comments, but we could also have used time between first and last post, or mean time between posts. We find that any of the three metrics shows roughly the same result, with more-active users more toxic on average (see Supplementary Section S.F). The number-of-comments method also has the advantage that each user contributes roughly the same number of comments to their bucket’s average.)
Figure 2: Trends in each metric over users’ lifetimes, from their first comment to their thousandth.
Shaded bands are 95% confidence intervals over all users in the relevant activity bucket. Trend lines are fit to the means over all users in an activity bucket.
The x-axis is logarithmic, for visibility of shorter activity buckets. For a linear x-axis scale and fit line, see Supplementary Section S.C.
A natural next question is whether the more active users made more toxic comments from the beginning, or if they became more toxic over their lifetime on the site.
@@
Individual plots have the same structure as Fig. 2, with 95% confidence intervals and trend lines fitted to bucket means.
We also see a consistent increase in toxicity score over users’ lifetimes. But knowing that our user sample spans over a decade of time, we might wonder if this effect was consistent over Reddit’s lifespan as well. In Figure 3, we separate users into cohorts based on the year in which they began commenting. This separation highlights one of the key findings from our analysis: in the earlier years, there is a downward trend for both toxicity and downvote scores, meaning that users become less toxic over time; but the trend gradually reverses over the years until, by 2022, there’s a clear upward slope, meaning that users become more toxic over time. (Patterns appear roughly the same whether the x-axis is logarithmic or linear - see Supplementary Section S.C.)
@@
The lower plots show the cumulative sum of the upper plots, in order to more clearly see when lines on the upper plot cross zero; that is, when the metric slopes become more positive than negative. The minimum value for each lifetime activity bucket is marked by a vertical line. This is the point where the slopes can be said to cross zero. Users who begin commenting after this point will tend to become more toxic over their lifetimes, rather than less.
For visibility, the slopes are scaled so that the maximum absolute value is 1 for each lifetime activity bucket.
A more granular view of these trends can be seen in Figure 4, where each point represents the slope of an entire lifetime trend line such as those in Figure 3. Here the trend is even more clear: lifetime toxicity slopes are negative in earlier years, and become positive in later years. When plotting them (upper plots), we see that the lines all cross 0 at around the same point. We can estimate the exact place where each line crosses zero using the cumulative sums (the lower plots); the minimums on these plots show the point where the values on the upper plots switch from being mostly negative to mostly positive.
@@
The structure is the same as Fig. 2, with 95% confidence intervals and trend lines fitted to bucket means.
Figure 5 sums up what we’ve learned so far. First, before the mid-2010s, users’ comments, on average, become less toxic the longer they’ve been on the site. Users who begin commenting after this point, however, make more toxic comments the longer they’ve been commenting. Second, more prolific users make more toxic comments than lower-activity users, though this effect is less strong when limiting to users from the same time period. Third, more prolific users score higher than others from their very first comment, rather than becoming toxic later in their lifetimes.
@@
Figure 6: Trends in toxicity over 6(a)) Wikipedia and 6(b)) Reddit users’ lifetimes, from their first comment to their thousandth.
6(b)) is a reproduction of Fig. 2 for ease of comparison.
@@
The structure is the same as Fig. 4. The lower plot shows the cumulative sum of the upper plot, in order to more clearly see when lines on the upper plot cross zero. The minimum value for each lifetime activity bucket is marked by a vertical line. This is the point where the slopes can be said to cross zero. For visibility, the slopes are scaled so that the maximum absolute value is 1 for each lifetime activity bucket.
Reddit, like most social media sites, has undergone many changes since its inception. To see if this was simply an artifact of culture change on a particular site, we can apply the same analysis to a second site: Wikipedia. While Wikipedia isn’t primarily a social media site, it has vigorous discussions between editors and is known for strongly enforced cultural norms.[44] As an analogy to Reddit comments, we use edits to Wikipedia “talk pages”, which are pages attached to Wikipedia articles or user accounts for discussion of edits to the relevant page. In Figure 6, we compare lifetime toxicity scores for Wikipedia users and Reddit users. The two are similar in many ways: there’s a clear ordering of users with different lifetime comment counts, and different groups on the same site show similar lifetime trends. But there’s also a striking difference: on Wikipedia, high-activity users are less toxic than less-active ones.
Figure 7 shows that there’s also a similar change from negative to positive slope in 2013. That is, as on Reddit, pre-2013 Wikipedia users became less toxic of their life cycles on the site, but post-2013 they became more toxic over their life cycles. (This is only apparent in the shortest-lived user buckets, but note how much smaller the other buckets are; the difference between bucket sizes is larger than on Reddit, especially when considering bucket sizes in the earlier history of the site, as in Supplementary Section S.G.)
It is interesting to ask how these findings about the contrast in toxicity by activity level — first for Reddit, then for Wikipedia — accord with intuition about social media. Earlier, it may have seemed obvious that the most active users were the most toxic, given social media’s reputation for toxic behavior. When Figure 6(a) shows the opposite, it might seem equally obvious that people with toxic behavior by definition don’t follow norms and can’t last long. But the point is that both of these “obvious” outcomes are observed, in different settings. To quote Duncan Watts in Everything is Obvious, “When every answer and its opposite appears equally obvious, then, as Lazarsfeld put it, ‘something is wrong with the entire argument of “obviousness".’[45]
The question is thus not which tendency is true of social media in general - whether high- or low-activity users are more toxic - but which, if any, is characteristic of a particular community. Viewed this way, as a comparison across different sites, it accords with some underlying intuition that high-activity Wikipedia users are less toxic than low-activity ones. Indeed, Wikipedia is known for its strongly enforced social norms, and is often favorably compared to other social media giants; Wikipedia has been described as “one of the few remaining places that retains the faintly utopian glow of the early World Wide Web”.[44]
@@
Shaded bands are 95% confidence intervals over the mean post scores for each time a URL was posted, calculated as the mean score for all comments on that post. Trend lines are fit to the means over all URLs in a bucket.
@@
The structure is the same as Fig. 2. Shaded bands are 95% confidence intervals over the score for the nth comment on each talk page. Trend lines are fit to the means over all pages in a bucket.
In the case of Reddit, we can follow specific URLs that are posted many times as top-level link posts. We focus on YouTube and Imgur URLs, since each points to to a fixed video or image, and both sites make it easy to detect different URLs that point to the same item. As with users, we order posts of the same URL in chronological order, from the first time the URL was posted to the last. When we apply a score intended to quantify toxicity, the score for each post is the mean score over all comments on the post.
@@
We’ve seen that similar patterns hold across two different metrics and two different datasets. Are these patterns specific to antisocial behavior, or would any metric of user behavior look similar? Are there metrics that don’t show the same patterns? In fact, every other metric we tried does look different from the patterns we’ve observed. In Supplementary Section S.E, we examine two kinds of scores for Reddit users: various binary thresholds on a comment’s upvote score, and “controversiality”, a binary measure defined by Reddit as comments with upvote/downvote ratios close to 0.5.[46] Recall the three main findings: users’ mean toxicity is correlated with their activity level; users change over their lifetimes; and the direction of change reverses in the mid-2010s.
For positive vote thresholds, there’s correlation between mean score and activity level, though the shape of user trajectories is clearly different (Fig S28). Lifetime slopes also don’t change much over time, though it’s interesting to note that lower-activity users have consistently negative trends, and higher-activity users consistently positive trends (Fig S32). For negative vote thresholds, there’s still some correlation between mean score and activity level, except for the least active users (Fig S29). However, different activity levels no longer have similar slopes, and except for the least extreme threshold, slopes don’t change over time (Fig S33). Controversiality, the metric that’s most orthogonal to our toxicity and downvote scores, also looks the most different from the patterns we found. It doesn’t show much of a difference between mean score for different activity levels (Fig S30). Slopes also show little difference between activity buckets and over time, and are generally positive (Fig S34).

### v2

@@ Abstract
An increasing amount of attention has been devoted to the problem of “toxic” or antisocial behavior on social media. In this paper we analyze such behavior at very large scales: over a 14-year time span on nearly 500 million comments from Reddit and Wikipedia, grounded in two different proxies for toxicity.
@@ Introduction
Antisocial behavior on social media is recognized as an increasingly important problem for the experience of users on many platforms. Often referred to with the umbrella term “toxicity”,[1] such behavior has been implicated in patterns of online harassment,[2, 3] outcomes of civic processes,[4, 5, 6, 7] and physical violence.[8, 9] However, its underlying mechanisms are still poorly understood. One reason for this challenge is that the culture and dynamics of social media are also constantly changing, both within and across platforms.[10, 11] Users also influence and are influenced by the communities they interact with. [12, 13]
@@
We note at the outset that it’s worth being critical of the term “toxicity”. The term is commonly assumed to need no definition,[1, 14] and often used without explanation - but some think it’s too vague to be meaningful, or even that it’s used to avoid more uncomfortable topics such as radicalization or identity-based hatred.[1, 15] It is therefore important to discuss the sense in which we will use the term in our work.
@@
Google’s Jigsaw, the makers of the Perspective API,[16, 17, 18] give a practical reason for choosing the term “toxicity”: when teaching humans to annotate their training data for generally problematic text, they found that “toxic” was the easiest concept for annotators to understand[19] That is, they tried various terms for various forms of problematic behavior, and the annotators were most likely to agree on which comments met the definition of “toxic” than for any of the other terms. For the API metric used in this work, “severe toxicity”, the specific definition they used was: “a hateful, aggressive, disrespectful comment or otherwise very likely to make a user leave a discussion or give up on sharing their perspective”.[20] Whether or not we agree with their definition,[21, 22, 23] the Perspective API model is clearly measuring some signal.[16] And since we’re interested in patterns at large scale over long time periods, we think it can give us useful information. Also, the fact that the API is commonly used in research[23] makes it easier to put results in context, and its use in real-world applications[24, 25] means that it’s worth studying.
@@
Another way to address the subtlety of the underlying concepts is to quantify them with multiple independent measures. Accordingly, we supplement the Perspective API score with an independent metric: a “downvote score” that also measures some kinds of antisocial behavior. Reddit’s content ranking depends on user voting, where any user can “upvote” or “downvote” a post or comment. Reddit asks users to vote based on whether a comment “contributes to conversation” (though in practice, users sometimes vote based on other factors, such as whether they agree with the comment.).[26] A comment’s score reflects community judgment, with negative scores (more downvotes than upvotes) often indicating toxic content that users found objectionable. While comments might also be downvoted for being unpopular, off-topic, or misleading, a negative score generally signals content the community deemed inappropriate or undesirable for wider visibility. As we will see, the broad conclusions we draw hold under both numerical metrics, despite their reliance on disjoint sources of information in the underlying data.
A large body of research addresses online toxicity and antisocial behavior, many specifically using the Perspective API toxicity score.[27, 28, 29, 30] Xia et al[27] studied various predictors of comment toxicity on Reddit, including toxicity of the author’s previous comments and of the previous comments in the discussion, while Avalle et al[28] found that comments increase in toxicity the longer a conversation lasts.
Other metrics of “toxicity” or “antisocial behavior” include banned users,[31] hand-coded “civility”,[32, 33] and custom-designed toxicity or hate speech classifiers.[34, 35] However, most existing work focuses on conversation-level[28, 27, 35, 36, 32, 37, 38] or community-level[39, 40, 41] patterns. Mamakos and Finkel[30] do track individual users between subreddits (communities within Reddit), finding that users who comment in politically partisan contexts are more toxic than users who don’t. Meanwhile, Chandrasekharan et al[42] follow changes in user population and overall toxicity over time for subreddits that were affected by quarantining. Sun et al[43] find that mean “incivility” across Reddit is fairly stable from 2006-2019, which accords with our findings on toxicity (Fig. S37).
@@
Few works follow individual user trajectories over time, as this work does. Of those that do, Cheng et al[31] analyze users who will go on to be banned from a community. Similar to this work, they find that “future banned users” are different from “never-banned users” from their very first post, and that these users’ behavior changes over time according to their main metrics: post deletion rate and text quality. They also measure slopes in users’ post deletion rates over time, though they focus on whether the slopes change between the first and second half of a user’s lifetime. Mekacher et al[44] track user toxicity over the first 150 days of users’ lifetimes on Gettr, a Twitter alternative popular with the far right, and find that lifetime toxicity is flat; however, age in days is a different measure of lifetime than the number-of-posts measure we use. Almerekhi et al[34] follow user change at a more local level, using a binary definition of “toxicity” and counting events where a user made a “toxic” comment after a “non-toxic” comment or vice versa. Waller and Anderson[39] also find a temporal shift in other characteristics at a similar time to the shift found in this work. They found a sharp change in user polarization, measured by the partisanship of communities that users commented in. However, this polarization event wasn’t relative to user lifetimes; existing users all polarized at approximately the same time, though most of the polarization was driven by new users.
The present work is interested in whether patterns of toxicity change over time, not just globally across a platform, but within individual users. We analyze trajectories across four dimensions: activity level (total lifetime number of comments), two different metrics of antisocial behavior (Perspective API toxicity score and “downvote score”, discussed below), two different social media sites (Reddit and Wikipedia), and two types of histories (people/users and information/URLs).
In nearly all cases, we observe a striking reversal: users do change over their lifetimes, but the direction of that change hasn’t been stable over time. Before the mid-2010s, the longer a user spent on a social media site, the less toxic they became. After the mid-2010s, the longer they spent on the site, the more toxic they became.
We also observe another pattern: First, low-activity users are generally separable from higher-activity users in their toxicity behavior. On Reddit, higher-activity users have higher toxicity scores; on Wikipedia, the reverse is true. This raises the question of how this difference comes about: one possibility is that high-activity users start out looking exactly like low-activity users and gradually reach their final level of toxicity; a second possibility is that high and low-activity users have different toxicity from the beginning. We find that the data is more consistent with the latter principle: Low- and high-activity users have different average toxicity metrics from their very first post. That is, it doesn’t seem that users start out the same, then only become more or less toxic the more they comment. Instead, they’re distinct from each other from the beginning.
@@ Results
The toxicity score is from the Perspective API,[45] a machine learning model that assigns a piece of text a score between 0 and 1, where 1 is most toxic. The API defines “toxicity” as “a very hateful, aggressive, disrespectful comment or otherwise very likely to make a user leave a discussion or give up on sharing their perspective”.[20] Scores for comments not in English were excluded, to prevent bias between models for different languages.[46]
@@
The downvote score comes from the Reddit voting system, where any user can “upvote” or “downvote” a comment based on whether it “contributes to conversation”.[26] Here “downvote score” for a group of comments denotes the percentage of comments that were downvoted, meaning comments that received more downvotes than upvotes.
@@
See Methods and Table 1 for more details. The Reddit user dataset contains all comments made by a random sample of 1 million users ($\sim 1\%$), for a total of 123 million comments. The Wikipedia dataset contains all edits made to English Wikipedia talk pages, after filtering out bot accounts, for a total of 122 million edits from 2.8 million accounts, on 21.8 million talk pages. Supplementary Section S.H shows the number of accounts created in different years.
The Reddit URL dataset consists of all Reddit posts of YouTube or Imgur URLs. Scores for each post were calculated as the mean of scores for all comments on the post. Downvote score for each post was defined as the percentage of comments on that post that were downvoted.
Black bars represent 95% confidence intervals over the mean, from 1,000 bootstrap runs.
@@
As a first question, we consider whether users who comment often are different from more infrequent users in the metrics we study (Fig 1). On average, these groups are indeed different, with the most prolific users showing the highest average for both toxicity and downvote scores. For toxicity scores, there is a clear upward trend as we move from less-active to more-active users.
For the downvote score, the most active users still have the highest average score, but here the trend is more U-shaped: moderate-activity users have the lowest scores, with the most- and least-active users both having high scores. An intuitive explanation for the U-shape might be that downvotes capture a wider range of factors than toxicity, such as off-topic or unpopular opinions, and that inexperienced commenters are more likely to break community norms and thus be downvoted for reasons other than toxicity[12]. Or there might be a selection effect: users whose comments are downvoted might be more likely to stop commenting, and thus have a lower lifetime number of comments.[47] However, our current analysis doesn’t allow us to determine if either explanation is accurate.
(To note, there are a number of options for stratifying users by activity. We use the lifetime number of comments, but we could also have used time between first and last post, or mean time between posts. We find that any of the three options show similar results, with more-active users more toxic on average (see Supplementary Section S.G). The number-of-comments method also has the advantage that each user contributes a comparable number of comments to their bucket’s average.)
Figure 2: Trends in each metric over users’ lifetimes, from their first comment to their thousandth. Users are grouped by their “activity” (lifetime comment count), and users with higher activity tend to have higher average toxicity scores.
Trend lines are fit to the means over all users in an activity bucket. Shaded bands are 95% confidence intervals for the mean.
The x-axis is logarithmic, for visibility of shorter activity buckets. For a linear x-axis scale and fit line, see Supplementary Section S.D.
Asterisks represent confidence that the relevant slope is positive or negative, for toxicity score and downvote score respectively. (See Supplementary Section S.B.2 for p-values.)
@@
Individual plots have the same structure as Fig. 2, with 95% confidence intervals over bucket means and trend lines fitted to the means.
A natural next question is whether the most-active users made more toxic comments from the beginning, or if they became more toxic over their lifetime on the site. In fact, in Fig. 2 we see a consistent increase in toxicity score over users’ lifetimes (less so for downvote score). But knowing that our user sample spans over a decade of time, we might wonder if this effect was consistent over Reddit’s lifespan as well. In Figure 3, we separate users into cohorts based on the year in which they began commenting. This separation highlights one of the key findings from our analysis: in the earlier years, there is a downward trend for both toxicity and downvote scores, meaning that users become less toxic over time; but the trend gradually reverses over the years until, by 2022, there’s a clear upward slope, meaning that users become more toxic over time. This appears for both toxicity and downvote score, though the reversal happens a few years later for toxicity score.
We calculate statistical significance of slopes using a two-sided t-test, where significance means that the true slope of the data is significantly different from zero (Section S.B.2). Similarly, 95% confidence intervals such as those in Fig. 4 are calculated using the standard error of the fitted slope. (Patterns appear similar whether the x-axis is logarithmic or linear - see Section S.D).
@@
The approximate point where each line crosses zero is marked with a vertical line. This is the time point where slopes go from mostly-negative (users become less toxic over their lifetime) to mostly-positive (users become more toxic over their lifetime). (Details in Supplementary Section S.C.)
Shaded areas represent 95% confidence intervals around the fitted slope. For visibility, the values are scaled so that the maximum absolute value is 1 for each lifetime activity bucket.
A more granular view of these trends can be seen in Figure 4, where each point represents the slope of an entire lifetime trend line such as those in Figure 3. Here the trend is even more clear: for both toxicity and downvote score, lifetime toxicity slopes are negative in earlier years and become positive in later years. When plotting them, we see that the lines all cross 0 at around the same time point (estimated using cumulative sums - see Supplementary Section S.C).
@@
The structure is the same as Fig. 2, with 95% confidence intervals around bucket means and trend lines fitted to the means.
Asterisks represent confidence levels for each slope fit, for the earlier and later cohort respectively.
Figure 5 sums up what we’ve learned so far. First, before the mid-2010s, users’ comments, on average, become less toxic the longer they’ve been on the site. Users who begin commenting after this point, however, make more toxic comments the longer they’ve been commenting. Second, more prolific users, on average, make more toxic comments than lower-activity users, though this effect is less strong when limiting to users from the same time period. Third, more prolific users score higher than others from their very first comment, rather than becoming toxic later in their lifetimes.
@@
Figure 6: Trends in toxicity over 6(a)) Wikipedia and 6(b)) Reddit users’ lifetimes, from their first comment to their thousandth. Unlike Reddit users, Wikipedia users with longer comment histories make comments that are less toxic, on average, than less-active users’ comments.
Panel 6(b)) is a reproduction of Fig. 2 for ease of comparison.
@@
The structure is the same as Fig. 4. Trends are much noisier than in Fig. 4 (note that most user-activity buckets have much smaller user counts than in the Reddit dataset.) The approximate point where each line crosses zero is marked with a vertical line. This is the point where slopes go from mostly-negative to mostly-positive. (Crossing points for this figure can be seen more clearly in Fig. S16.)
Reddit, like most social media sites, has undergone many changes since its inception. To see if this was simply an artifact of culture change on a particular site, we can apply the same analysis to a second site: Wikipedia. While Wikipedia isn’t primarily a social media site, it hosts vigorous discussions between editors and is known for strongly enforced cultural norms.[48] As an analogy to Reddit comments, we use edits to Wikipedia “talk pages”, which are pages attached to Wikipedia articles or user accounts for discussion of edits to the relevant page. In Figure 6, we compare lifetime toxicity scores for Wikipedia users and Reddit users. The two are similar in many ways: there’s a clear ordering of users with different lifetime comment counts, and different groups on the same site show similar lifetime trends. But there’s also a striking difference: on Wikipedia, high-activity users are less toxic than less-active ones.
Figure 7 shows that there’s also a similar change from negative to positive slope in 2013. That is, as on Reddit, pre-2013 Wikipedia users became less toxic over their life cycles on the site, but post-2013 users became more toxic over their life cycles. (This is only apparent in the shortest-lived user buckets, but note how much smaller the other buckets are; the difference between bucket sizes is larger than on Reddit, especially when considering bucket sizes in the earlier history of the site, as in Supplementary Section S.H.)
@@
Shaded bands are 95% confidence intervals over the post scores for each time a URL was posted. Post scores are calculated as the mean score for all comments on that post. Trend lines are fit to the mean post score over all URLs in a bucket.
@@
Asterisks represent confidence that the relevant slope is positive or negative, for toxicity score and downvote score respectively.
@@
The structure is the same as Fig. 2. Shaded bands are 95% confidence intervals over the mean score for the nth comment on each talk page. Trend lines are fit to the means over all pages in a bucket.
In the case of Reddit, we can follow specific URLs that are posted many times as top-level link posts. We focus on YouTube and Imgur URLs, since each points to a fixed video or image, and both sites make it easy to detect different URLs that point to the same item. As with users, we order posts of the same URL in chronological order, from the first time the URL was posted to the last. When we apply a score intended to quantify toxicity, the score for each post is the mean score over all comments on the post.
@@
We’ve seen that similar patterns hold across two different metrics and two different datasets. Are these patterns specific to antisocial behavior, or would any metric of user behavior look similar? In fact, every metric we tried, except for those related to antisocial behavior, does look different from the patterns we’ve observed. In Supplementary Section S.F, we examine two kinds of scores for Reddit users: various binary thresholds on a comment’s upvote score, and “controversiality”, a binary measure defined by Reddit as comments with upvote/downvote ratios close to 1.[49]
Recall the three main findings: users’ mean toxicity is correlated with their activity level; users change over their lifetimes; and the direction of change reverses in the mid-2010s. For positive vote thresholds, there’s correlation between mean score and activity level, though the shape of user trajectories is clearly different (Fig S39). Lifetime slopes also don’t change much over time, though it’s interesting to note that lower-activity users have consistently negative trends, and higher-activity users have consistently positive trends (Fig S43). For negative vote thresholds, there’s still some correlation between mean score and activity level, except for the least active users (Fig S40). However, different activity levels no longer have similar slopes, and for the more extreme thresholds, slopes don’t change over time (Fig S44). Controversiality, the metric that’s most orthogonal to our toxicity and downvote scores, also looks the most different from the patterns we found. Controversiality doesn’t show much of a difference between mean score for different activity levels (Fig S41). Slopes also show little difference between activity buckets and over time, and are generally positive (Fig S45).